# Newton-Flow Particle Filters based on Generalized Cramér Distance

### v1

@@ VI-B Gradient
$=$ $4\,w_{k}\Bigg{\{}\sum_{\begin{subarray}{c}i=1\\ i\neq k\end{subarray}}^{L}w_{i}\,(\underline{x}_{k}-\underline{x}_{i})\log\big{(}\|\underline{x}_{k}-\underline{x}_{i}\|_{2}^{2}\big{)}$
$\quad\quad-\sum_{j=1}^{M}\tilde{w}_{j}\,(\underline{x}_{k}-\tilde{\underline{x}}_{j})\log\big{(}\|\underline{x}_{k}-\tilde{\underline{x}}_{i}\|_{2}^{2}\big{)}\Bigg{\}}$
$+K_{2}\,w_{k}\Bigg{\{}\sum_{i=1}^{L}w_{i}\,\underline{x}_{i}-\sum_{j=1}^{M}\tilde{w}_{j}\,\tilde{\underline{x}}_{j}\Bigg{\}}$
$K_{2}=2(K_{1}-2)$ is a large constant penalizing the difference between the means of $\tilde{f}_{e}(\underline{x})$ and $f_{e}(\underline{x})$.
@@ VI-C Hesse Matrix
${\mathbf{H}}=\begin{bmatrix}{\mathbf{H}}_{11}&{\mathbf{H}}_{12}&{\mathbf{H}}_{13}&\cdots&{\mathbf{H}}_{1L}\\ {\mathbf{H}}_{21}&{\mathbf{H}}_{22}&{\mathbf{H}}_{23}&\cdots&{\mathbf{H}}_{2L}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ {\mathbf{H}}_{L1}&{\mathbf{H}}_{L2}&{\mathbf{H}}_{L3}&\cdots&{\mathbf{H}}_{LL}\end{bmatrix}\enspace$. (55)
@@
for $k=1,\ldots,L$. The off-diagonal block matrices are given by
@@
for $k=1,\ldots,L$, $l=1,\ldots,L$, $l\neq k$. The Hesse matrix is symmetric.

### v2

@@ VI-B Gradient
The gradient of $D$ in (48) w.r.t. $\underline{x}_{k}$ is given by [^4] For minimization, we use the squared distance $D$ for simplicity.
@@
$=$ $4\,w_{k}\Bigg\{\sum_{\begin{subarray}{c}i=1\\ i\neq k\end{subarray}}^{L}w_{i}\,\underline{\mathcal{G}}(\underline{x}_{k}-\underline{x}_{i})-\sum_{j=1}^{M}\tilde{w}_{j}\,\underline{\mathcal{G}}(\underline{x}_{k}-\tilde{\underline{x}}_{j})\Bigg\}$
$+2\,K\,w_{k}\Bigg\{\sum_{i=1}^{L}w_{i}\,\underline{x}_{i}-\sum_{j=1}^{M}\tilde{w}_{j}\,\tilde{\underline{x}}_{j}\Bigg\}$
with
$\underline{\mathcal{G}}(\underline{d})=\left[\frac{\underline{d}^{T}\underline{d}}{d_{\text{min}}^{2}+\underline{d}^{T}\underline{d}}+\log\bigl(d_{\text{min}}^{2}+\underline{d}^{T}\underline{d}\bigr)\right]\,\underline{d}\kern 5.0pt.$ (55)
@@ VI-C Hesse Matrix
${\mathbf{H}}=\begin{bmatrix}{\mathbf{H}}_{11}&{\mathbf{H}}_{12}&{\mathbf{H}}_{13}&\cdots&{\mathbf{H}}_{1L}\\ {\mathbf{H}}_{21}&{\mathbf{H}}_{22}&{\mathbf{H}}_{23}&\cdots&{\mathbf{H}}_{2L}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ {\mathbf{H}}_{L1}&{\mathbf{H}}_{L2}&{\mathbf{H}}_{L3}&\cdots&{\mathbf{H}}_{LL}\end{bmatrix}\kern 5.0pt$. (56)
@@
${\mathbf{H}}_{kk}$ $=\frac{\partial G_{k}}{\partial\underline{x}_{k}^{T}}=\frac{\partial^{2}D}{\partial\underline{x}_{k}\partial\underline{x}_{k}^{T}}$ (57)
$=2\,K\,w_{k}^{2}\,{\mathbf{I}}_{N}$
@@
for $k=1,\ldots,L$ with
@@
$+\left[\frac{\underline{d}^{T}\underline{d}}{d_{\text{min}}^{2}+\underline{d}^{T}\underline{d}}+\log\bigl(d_{\text{min}}^{2}+\underline{d}^{T}\underline{d}\bigr)\right]\,{\mathbf{I}}_{N}\kern 5.0pt,$
where ${\mathbf{I}}_{N}$ is the $N$-dimensional identity matrix. The off-diagonal block matrices are given by
${\mathbf{H}}_{kl}$ $=\frac{\partial G_{k}}{\partial\underline{x}_{l}^{T}}=\frac{\partial^{2}D}{\partial\underline{x}_{k}\partial\underline{x}_{l}^{T}}$ (59)
[…]
for $k=1,\ldots,L$, $l=1,\ldots,L$, $l\neq k$. The Hesse matrix is symmetric and positive definite.